# Comprehensive Review of Neural Differential Equations for Time Series Analysis

### v1

@@ Abstract
Time series modeling and analysis has become critical in various domains. Conventional methods such as RNNs and Transformers, while effective for discrete-time and regularly sampled data, face significant challenges in capturing the continuous dynamics and irregular sampling patterns inherent in real-world scenarios. Neural Differential Equations (NDEs) represent a paradigm shift by combining the flexibility of neural networks with the mathematical rigor of differential equations. This paper presents a comprehensive review of NDE-based methods for time series analysis, including neural ordinary differential equations, neural controlled differential equations, and neural stochastic differential equations. We provide a detailed discussion of their mathematical formulations, numerical methods, and applications, highlighting their ability to model continuous-time dynamics. Furthermore, we address key challenges and future research directions. This survey serves as a foundation for researchers and practitioners seeking to leverage NDEs for advanced time series analysis.
@@ 1 Introduction
The exponential growth of time series data across diverse domains has necessitated more sophisticated analytical approaches, because of the complex nonlinear nature, irregular sampling, missing values, and continuous latent dynamics. Traditional approaches (Durbin and Koopman, 2012; Box et al., 2015) assume regular sampling and often linear relationships, limiting their applicability to real-world scenarios. While deep learning methods, including Recurrent Neural Network (RNN) (Rumelhart et al., 1986; Medsker and Jain, 1999), Long Short-term Memory (LSTM) (Hochreiter and Schmidhuber, 1997), Gated Recurrent Unit (GRU) (Chung et al., 2014), and Transformer (Vaswani et al., 2017) have shown promise in capturing nonlinearity, they remain constrained by discrete-time formulations (Che et al., 2018; Sun et al., 2020; Weerakody et al., 2021)
Neural Differential Equations (NDEs) emerged as a paradigm shift, offering a principled framework for continuous-time modeling with neural networks. The seminal work on Neural Ordinary Differential Equations (NODEs) (Chen et al., 2018) introduced continuous-time hidden state evolution, spawning numerous advances including Neural Controlled Differential Equations (NCDEs) (Kidger et al., 2020) incorporated external control paths, and Neural Stochastic Differential Equations (NSDEs) (Han et al., 2017; Tzen and Raginsky, 2019; Li et al., 2020; Oh et al., 2024) model uncertainty through stochasticity.
@@ 2 Preliminaries
Time series modeling seeks to represent sequential data ${\bm{x}}=(x_{0},x_{1},\ldots,x_{n})$, where each $x_{i}\in\mathbb{R}^{d_{x}}$, as a continuous latent process ${\bm{z}}(t)\in\mathbb{R}^{d_{z}}$ over a time domain $[0,T]$. As shown Figure 1, while interpolation methods simply fit approximation through observations (black dots), and RNNs with zero-order hold operate in discrete time steps, NDEs learn the actual dynamics by neural network $f(t)\equiv{{\mathrm{d}}{\bm{z}}(t)}/{{\mathrm{d}}t}$ that generated the data.
@@
Figure 1: Conceptual comparison of time series modeling methods. Piecewise interpolation and traditional RNNs update at discrete points, approximating segments of trajectory. NDE-baed methods learn continuous temporal dynamics, capturing smooth and data-driven evolution.
By parameterizing the temporal evolution of ${\bm{z}}(t)$ with neural networks, NDEs provide a flexible framework for handling irregularly sampled data, missing observations, and long-horizon dynamics, often with more memory efficiency than standard discrete-time architectures (Chen et al., 2018; Rubanova et al., 2019; Kidger et al., 2020; Oh et al., 2024).
RNNs, including LSTMs (Hochreiter and Schmidhuber, 1997) and GRUs (Chung et al., 2014), operate in discrete time, updating hidden states sequentially based on observed inputs. Transformer-based models (Vaswani et al., 2017) offer an alternative by leveraging self-attention mechanisms, though they typically rely on fixed positional encodings and do not inherently capture continuous-time dynamics.
Various modification have been proposed to introduce a notion of time awareness in RNNs (Che et al., 2018; Cao et al., 2018; Rajkomar et al., 2018). One approach adjusts updates for time intervals, while another uses decay mechanisms to approximate continuous evolution. However, these heuristics do not always generalize across different settings and, in some cases, have been found to perform comparably to standard RNNs (Mozer et al., 2017; Rubanova et al., 2019).
@@ 2.1 Neural Ordinary Differential Equations
NODEs (Chen et al., 2018) model the latent state ${\bm{z}}(t)$ as:
@@
where ${\bm{z}}(0)=h({\bm{x}};\theta_{h})$, $h:\mathbb{R}^{d_{x}}\to\mathbb{R}^{d_{z}}$ is a neural network parameterized by $\theta_{h}$, and $f\bigl{(}s,{\bm{z}}(s);\theta_{f}\bigr{)}$ approximates $\frac{\mathrm{d}{\bm{z}}(t)}{\mathrm{d}t}$. The vector field $f$ is typically implemented using multi-layer perceptrons or more sophisticated architectures. One can view these approaches as an infinite-layer generalization of residual networks, allowing integration over time rather than discrete stacking of layers (Chen et al., 2018; Rubanova et al., 2019; Dupont et al., 2019; Kidger et al., 2020).
@@ 2.2 Neural Controlled Differential Equations
NCDEs (Kidger et al., 2020) extend NODEs by incorporating a control path $X(t)$ for updating state over time:
@@
where the integral is interpreted as a Riemann–Stieltjes integral, allowing for discontinuous paths. Piecewise-smooth control paths $X(t)$ are typically constructed using natural cubic splines (Kidger et al., 2020) or Hermite cubic splines (Morrill et al., 2021a). Furthermore, Morrill et al. (2021b) extend Neural CDEs by incorporating rough path theory for the generalized formulation.
@@ 2.3 Neural Stochastic Differential Equations
NSDEs (Han et al., 2017; Tzen and Raginsky, 2019; Li et al., 2020) incorporate stochasticity through:
${\bm{z}}(t)={\bm{z}}(0)+\int_{0}^{t}f\bigl{(}s,{\bm{z}}(s);\theta_{f}\bigr{)}\,\mathrm{d}s+\int_{0}^{t}g\bigl{(}s,{\bm{z}}(s);\theta_{g}\bigr{)}\,\mathrm{d}W(s),$ (3)
where $W(t)$ is a Wiener process (or Brownian motion), $f\bigl{(}\cdot;\theta_{f}\bigr{)}$ is the drift function, and $g\bigl{(}\cdot;\theta_{g}\bigr{)}$ is the diffusion function. The stochastic integral follows Itô or Stratonovich interpretation. While NODEs describe deterministic evolution through ordinary differential equations, NSDEs model uncertainty and noise through Brownian motion terms (Jia and Benson, 2019; Liu et al., 2019). This extension enables more robust modeling of real-world phenomena where randomness plays a crucial role (Kidger et al., 2021b, c; Oh et al., 2024), each offering different approaches to balancing stability, expressivity, and computational tractability.
@@ 3 Theoretical Considerations
The theoretical foundations of NDEs are essential for understanding their ability to model complex temporal dynamics. Their connection to dynamical systems explains how they capture continuous-time evolution, while universal approximation properties justify their flexibility in learning intricate patterns. Additionally, stability and convergence analysis ensures robustness, making NDEs a reliable framework.
@@ 3.1 Dynamical Systems Perspective
NDEs generalize the notion of discrete layers (as in RNNs or Transformers) to a continuous-time vector field. According to equation (1), the latent representation ${\bm{z}}(t)$ evolves:
@@
Here, $X(t)$ is typically a spline or piecewise interpolation of observed data ${\bm{x}}$, as explained in equation (2).
Stochastic extensions, such as NSDEs, incorporate Brownian motion $W(t)$ or jump processes to model noise or uncertainty in the underlying dynamics, as shown equation (3):
$\mathrm{d}{\bm{z}}(t)=f\bigl{(}{\bm{z}}(t);\theta_{f}\bigr{)}\,\mathrm{d}t+g\bigl{(}{\bm{z}}(t);\theta_{g}\bigr{)}\,\mathrm{d}W(t),$ (6)
@@
In parallel with NDEs, explicit approaches like Neural Flows have also been proposed (Lu et al., 2018; Sonoda and Murata, 2019; Grathwohl et al., 2019; Massaroli et al., 2020a; Biloš et al., 2021). Rather than parameterizing the rate of change $\mathrm{d}{\bm{z}}(t)/\mathrm{d}t$, Neural Flows directly model the solution map:
@@
where ${\bm{z}}(0)=h({\bm{x}};\theta_{h})$ is the initial condition. This explicit representation bypasses the need for numerical ODE solvers, potentially offering faster computations or enhanced stability via invertibility constraints (Kobyzev et al., 2020; Zhang et al., 2020; Papamakarios et al., 2021). Such architectures align with normalizing-flow models, which are invertible by construction and allow for efficient likelihood evaluation.
@@
The distinction between NODEs/NCDEs (which define $\mathrm{d}{\bm{z}}/\mathrm{d}t$ and rely on solvers) and Neural Flows (which specify ${\bm{z}}(t)$ directly) parallels the classic implicit versus explicit solution dichotomy in differential equations. Implicit, solver-based methods (like NODEs, NCDEs, NSDEs) can handle irregular inputs through fixed or adaptive step sizes and are more general in capturing unknown or complex dynamics. Explicit approaches, like Neural Flows (Biloš et al., 2021), may yield simpler training pipelines for stable transformations, yet can suffer from constraints on network forms and from instabilities if the sought-after closed-form solution is nonexistent or approximate (Oh et al., 2025).
@@ 3.2 Universal Approximation Properties
Chen et al. (2018) established that continuous-time models can be viewed as an infinite-depth limit of residual networks, enabling them to approximate diffeomorphic transformations under mild regularity conditions, as represented in equation (4). However, autonomous ODE flows may limit expressivity by constraining trajectories to remain diffeomorphic transformations of initial states, thus prohibiting intersecting paths in latent space. Dupont et al. (2019) addressed these limitations through Augmented ODEs, which append an auxiliary variable $\bm{a}(t)$ to the state:
$\frac{\mathrm{d}}{\mathrm{d}t}\begin{pmatrix}{\bm{z}}(t)\\ \bm{a}(t)\end{pmatrix}=f\Bigl{(}{\bm{z}}(t),\bm{a}(t);\theta_{f}\Bigr{)}.$ (9)
@@
While NODEs rely solely on a learned vector field and initial value, Kidger et al. (2020) showed that Neural CDEs embed a control path $X(t)$ into the model as explained in equation (5): By allowing data updates to arrive as increments $\mathrm{d}X(t)$ at arbitrary times, NCDEs implement a form of continuous recursion, demonstrating universal approximation for continuous paths under sufficiently rich control signals. Extensions to rough paths (Morrill et al., 2021b) further expand coverage to long, irregular time series, handling signals too irregular for traditional ODE integrators.
@@
Stochastic extensions achieve universal approximation for broad classes of continuous-time random processes by learning drift and diffusion terms. Han et al. (2017) and Tzen and Raginsky (2019) formalized how parameterizing $f$ and $g$ via neural networks lets these models approximate various stochastic phenomena, such as Geometric Brownian motion or Ornstein–Uhlenbeck processes. Subsequent works generalize to jump processes and specialized SDE families (Jia and Benson, 2019; Li et al., 2020; Oh et al., 2024; Zhang et al., 2024), underscoring the versatility of NSDEs in capturing noise-driven dynamics.
@@ 3.3 Existence and Uniqueness of Solutions
For NODEs as explained in equation (4), classic results such as the Picard–Lindelöf theorem guarantee that a unique solution ${\bm{z}}(t)$ exists if the vector field $f$ is Lipschitz continuous in ${\bm{z}}$. Formally, if
@@
for some constant $L>0$, then the integral equation describing ${\bm{z}}(t)$ has a unique solution starting from ${\bm{z}}(0)=h({\bm{x}})$ (Chen et al., 2018). In practice, spectral normalization or weight clipping can enforce such Lipschitz constraints, improving both existence properties and training stability (Massaroli et al., 2020b; Rackauckas et al., 2020).
In NCDEs as shown equation (5), existence and uniqueness in this setting hinge on Lipschitz-like conditions for $f$ and on the path $X(t)$ possessing sufficient regularity (e.g., bounded variation or spline-based interpolation) (Kidger et al., 2020). If $X(t)$ is discontinuous or highly irregular, rough path theory provides a generalized framework under which well-posedness can still be established (Morrill et al., 2021a).
@@
NSDEs embed stochasticity via equation (6), existence and uniqueness then require both $f$ and $g$ to satisfy Lipschitz and linear-growth conditions in ${\bm{z}}$ (Tzen and Raginsky, 2019; Oh et al., 2024). One common conditions for the existence and uniquess of ${\bm{z}}$ are Lipschitz continuity and linear growth conditions for $f$ and $g$ such that if there exist constants $L$ and $C$ such that
$\|f\bigl{(}{\bm{z}}(t_{1});\theta_{f}\bigr{)}-f\bigl{(}{\bm{z}}(t_{2});\theta_{f}\bigr{)}\|+\|g\bigl{(}{\bm{z}}(t_{1});\theta_{g}\bigr{)}-g\bigl{(}{\bm{z}}(t_{2});\theta_{g}\bigr{)}\|$ (11)
$\,\leq\,L\|{\bm{z}}(t_{1})-{\bm{z}}(t_{2})\|,$
@@
then a unique strong solution ${\bm{z}}(t)$ exists for all finite $t$. Under these conditions, finite-time blow-up is prevented, ensuring the SDE is well-posed (Jia and Benson, 2019; Li et al., 2020). Langevin-type SDEs, linear-noise SDEs, and geometric SDEs—variants of Neural SDEs—are also designed to satisfy these conditions, thereby ensuring the existence and uniqueness of their solutions (Oh et al., 2024).

### v2

@@ Abstract
Time series modeling and analysis have become critical in various domains. Conventional methods such as RNNs and Transformers, while effective for discrete-time and regularly sampled data, face significant challenges in capturing the continuous dynamics and irregular sampling patterns inherent in real-world scenarios. Neural Differential Equations (NDEs) represent a paradigm shift by combining the flexibility of neural networks with the mathematical rigor of differential equations. This paper presents a comprehensive review of NDE-based methods for time series analysis, including neural ordinary differential equations, neural controlled differential equations, and neural stochastic differential equations. We provide a detailed discussion of their mathematical formulations, numerical methods, and applications, highlighting their ability to model continuous-time dynamics. Furthermore, we address key challenges and future research directions. This survey serves as a foundation for researchers and practitioners seeking to leverage NDEs for advanced time series analysis.
@@ 1 Introduction
The exponential growth of time series data across diverse domains has necessitated more sophisticated analytical approaches, because of the complex nonlinear nature, irregular sampling, missing values, and continuous latent dynamics. Traditional approaches Durbin and Koopman (2012); Box et al. (2015) assume regular sampling and often linear relationships, limiting their applicability to real-world scenarios. While deep learning methods, including Recurrent Neural Network (RNN) Rumelhart et al. (1986); Medsker and Jain (1999), Long Short-term Memory (LSTM) Hochreiter and Schmidhuber (1997), Gated Recurrent Unit (GRU) Chung et al. (2014), and Transformer Vaswani et al. (2017) have shown promise in capturing nonlinearity, they remain constrained by discrete-time formulations Che et al. (2018); Sun et al. (2020); Weerakody et al. (2021)
Neural Differential Equations (NDEs) emerged as a paradigm shift, offering a principled framework for continuous-time modeling with neural networks. The seminal work on Neural Ordinary Differential Equations (NODEs) Chen et al. (2018) introduced continuous-time hidden state evolution, spawning numerous advances including Neural Controlled Differential Equations (NCDEs) Kidger et al. (2020) incorporating external control paths, and Neural Stochastic Differential Equations (NSDEs) Han et al. (2017); Tzen and Raginsky (2019); Li et al. (2020); Oh et al. (2024) modeling uncertainty through stochasticity.
@@ 2 Preliminaries
Time series modeling seeks to represent sequential data ${\bm{x}}=(x_{0},x_{1},\ldots,x_{n})$, where each $x_{i}\in\mathbb{R}^{d_{x}}$, as a continuous latent process ${\bm{z}}(t)\in\mathbb{R}^{d_{z}}$ over a time domain $[0,T]$. Figure 1 illustrates the conceptual differences between conventional time series modeling and NDE-based methods. While interpolation methods simply fit curves through observations (black dots) and RNNs (often using zero-order hold) operate in discrete time steps, NDEs aim to learn the underlying continuous dynamics, ${\mathrm{d}}{\bm{x}}(t)/{\mathrm{d}}t$ (or ${\mathrm{d}}{\bm{z}}(t)/{\mathrm{d}}t$), parameterized by a neural network Chen et al. (2018). This parameterization provides NDEs with a flexible framework for handling challenges common in real-world time series, such as irregular sampling, missing observations, and capturing long-horizon dependencies. Furthermore, NDEs can offer greater memory efficiency compared to standard discrete-time architectures Chen et al. (2018); Kidger et al. (2020); Rubanova et al. (2019); Oh et al. (2024).
@@
Figure 1: Conceptual comparison of time series modeling. (a) Piecewise interpolation provides continuity but lacks dynamic modeling capabilities. (b) Traditional RNNs operate on discrete steps, encountering limitations with continuous dynamics, irregularly sampled data, and memory efficiency for long sequences. (c) NDE-based methods, such as NODEs, NCDEs, and NSDEs, learn continuous dynamics, inherently accommodate irregular sampling, and offer potential memory advantages, when employing techniques like the adjoint method.
RNNs, including LSTMs Hochreiter and Schmidhuber (1997) and GRUs Chung et al. (2014), operate in discrete time, updating hidden states sequentially based on observed inputs. Transformer-based models Vaswani et al. (2017) offer an alternative by leveraging self-attention mechanisms, though they typically rely on fixed positional encodings and do not inherently capture continuous-time dynamics.
Various modifications have been proposed to introduce a notion of time awareness in RNNs Che et al. (2018); Rajkomar et al. (2018). One approach adjusts updates for time intervals, while another uses decay mechanisms to approximate continuous evolution. However, these heuristics do not always generalize across different settings and, in some cases, have been found to perform comparably to standard RNNs Mozer et al. (2017); Rubanova et al. (2019).
@@ 2.1 Neural Ordinary Differential Equations
NODEs Chen et al. (2018) model the latent state ${\bm{z}}(t)$ as:
@@
where ${\bm{z}}(0)=h({\bm{x}};\theta_{h})$, $h:\mathbb{R}^{d_{x}}\to\mathbb{R}^{d_{z}}$ is a neural network parameterized by $\theta_{h}$, and $f\bigl{(}t,{\bm{z}}(t);\theta_{f}\bigr{)}$ approximates $\frac{\mathrm{d}{\bm{z}}(t)}{\mathrm{d}t}$. The vector field $f$ is typically implemented using multi-layer perceptrons or more sophisticated architectures. One can view these approaches as an infinite-layer generalization of residual networks, allowing integration over time rather than discrete stacking of layers Chen et al. (2018); Rubanova et al. (2019); Dupont et al. (2019); Kidger et al. (2020).
@@ 2.2 Neural Controlled Differential Equations
NCDEs Kidger et al. (2020) extend NODEs by incorporating a control path $X(t)$ for updating state over time:
@@
where the integral is interpreted as a Riemann–Stieltjes integral, allowing for discontinuous paths. Piecewise-smooth control paths $X(t)$ are typically constructed using natural cubic splines Kidger et al. (2020) or Hermite cubic splines Morrill et al. (2021a). Furthermore, Morrill et al. (2021b) extend Neural CDEs by incorporating rough path theory for the generalized formulation.
@@ 2.3 Neural Stochastic Differential Equations
NSDEs Han et al. (2017); Tzen and Raginsky (2019); Li et al. (2020) incorporate stochasticity through:
${\bm{z}}(t)={\bm{z}}(0)$ $+\int_{0}^{t}f\bigl{(}s,{\bm{z}}(s);\theta_{f}\bigr{)}\,\mathrm{d}s$
$+\int_{0}^{t}g\bigl{(}s,{\bm{z}}(s);\theta_{g}\bigr{)}\,\mathrm{d}W(s),$ (3)
where $W(t)$ is a Wiener process (or Brownian motion), $f\bigl{(}\cdot;\theta_{f}\bigr{)}$ is the drift function, and $g\bigl{(}\cdot;\theta_{g}\bigr{)}$ is the diffusion function. The stochastic integral follows Itô or Stratonovich interpretations. While NODEs describe deterministic evolution through ordinary differential equations, NSDEs model uncertainty and noise through Brownian motion terms Jia and Benson (2019); Liu et al. (2019). This extension enables more robust modeling of real-world phenomena where randomness plays a crucial role Kidger et al. (2021b, c); Oh et al. (2024), each offering different approaches to balance stability, expressivity, and computational tractability.
@@ 3.1 Dynamical Systems Perspective
NDEs generalize the notion of discrete layers (as in RNNs or Transformers) to a continuous-time vector field. According to Eq. (1), the latent representation ${\bm{z}}(t)$ evolves:
@@
Here, $X(t)$ is typically a spline or piecewise interpolation of observed input data ${\bm{x}}$, as explained in Eq. (2).
Stochastic extensions, such as NSDEs, incorporate Brownian motion $W(t)$ or jump processes to model noise or uncertainty in the underlying dynamics, as shown in Eq. (2.3):
$\mathrm{d}{\bm{z}}(t)=f\bigl{(}t,{\bm{z}}(t);\theta_{f}\bigr{)}\,\mathrm{d}t+g\bigl{(}t,{\bm{z}}(t);\theta_{g}\bigr{)}\,\mathrm{d}W(t),$ (6)
@@
In parallel with NDEs, explicit approaches like Neural Flows have also been proposed Lu et al. (2018); Sonoda and Murata (2019); Grathwohl et al. (2019); Massaroli et al. (2020a); Biloš et al. (2021). Rather than parameterizing the rate of change $\mathrm{d}{\bm{z}}(t)/\mathrm{d}t$, Neural Flows directly model the solution map:
@@
where ${\bm{z}}(0)=h({\bm{x}};\theta_{h})$ is the initial condition. This explicit representation bypasses the need for numerical ODE solvers, potentially offering faster computations or enhanced stability via invertibility constraints Kobyzev et al. (2020); Papamakarios et al. (2021). Such architectures align with normalizing-flow models, which are invertible by construction and allow for efficient likelihood evaluation.
@@
The distinction between NODEs/NCDEs (which define $\mathrm{d}{\bm{z}}/\mathrm{d}t$ and rely on solvers) and Neural Flows (which specify ${\bm{z}}(t)$ directly) parallels the classic implicit versus explicit solution dichotomy in differential equations. Implicit, solver-based methods (like NODEs, NCDEs, NSDEs) can handle irregular inputs through fixed or adaptive step sizes and are more general in capturing unknown or complex dynamics. Explicit approaches, like Neural Flows Biloš et al. (2021), may yield simpler training pipelines for stable transformations, yet can suffer from constraints on network forms and from instabilities if the sought-after closed-form solution is nonexistent or approximate Oh et al. (2025).
@@ 3.2 Universal Approximation Properties
Chen et al. (2018) established that continuous-time models can be viewed as an infinite-depth limit of residual networks, enabling them to approximate diffeomorphic transformations under mild regularity conditions, as represented in Eq. (4). However, autonomous ODE flows may limit expressivity by constraining trajectories to remain diffeomorphic transformations of initial states, thus prohibiting intersecting paths in latent space. Dupont et al. (2019) addressed these limitations through augmented ODEs, which append an auxiliary variable $\bm{a}(t)$ to the state:
$\frac{\mathrm{d}}{\mathrm{d}t}\begin{pmatrix}{\bm{z}}(t)\\ \bm{a}(t)\end{pmatrix}=f\Bigl{(}t,{\bm{z}}(t),\bm{a}(t);\theta_{f}\Bigr{)}.$ (9)
@@
While NODEs rely solely on a learned vector field and initial value, Kidger et al. (2020) showed that Neural CDEs embed a control path $X(t)$ into the model as explained in Eq. (5): By allowing data updates to arrive as increments $\mathrm{d}X(t)$ at arbitrary times, NCDEs implement a form of continuous recursion, demonstrating universal approximation for continuous paths under sufficiently rich control signals. Extensions to rough paths Morrill et al. (2021b) further expand coverage to long, irregular time series, handling signals too irregular for traditional ODE integrators.
@@
Stochastic extensions achieve universal approximation for broad classes of continuous-time random processes by learning drift and diffusion terms. Han et al. (2017) and Tzen and Raginsky (2019) formalized how parameterizing $f$ and $g$ via neural networks lets these models approximate various stochastic phenomena, such as Geometric Brownian motion or Ornstein–Uhlenbeck processes. Subsequent works generalize to jump processes and specialized SDE families Jia and Benson (2019); Li et al. (2020); Oh et al. (2024); Zhang et al. (2024), underscoring the versatility of NSDEs in capturing noise-driven dynamics.
@@ 3.3 Existence and Uniqueness of Solutions
For NODEs as explained in Eq. (4), classic results such as the Picard–Lindelöf theorem guarantees that a unique solution ${\bm{z}}(t)$ exists if the vector field $f$ is Lipschitz continuous in ${\bm{z}}$. Formally, if
@@
for some constant $L>0$, then the integral equation describing ${\bm{z}}(t)$ has a unique solution starting from ${\bm{z}}(0)=h({\bm{x}})$ Chen et al. (2018). In practice, spectral normalization or weight clipping can enforce such Lipschitz constraints, improving both existence properties and training stability Massaroli et al. (2020b); Rackauckas et al. (2020).
In NCDEs as shown in Eq. (5), existence and uniqueness in this setting hinge on Lipschitz-like conditions for $f$ and on the path $X(t)$ possessing sufficient regularity (e.g., bounded variation or spline-based interpolation) Kidger et al. (2020). If $X(t)$ is discontinuous or highly irregular, rough path theory provides a generalized framework under which well-posedness can still be established Morrill et al. (2021a).
@@
NSDEs embed stochasticity via Eq. (6), existence and uniqueness then require both $f$ and $g$ to satisfy Lipschitz and linear-growth conditions in ${\bm{z}}$ Tzen and Raginsky (2019); Oh et al. (2024). Common conditions for the existence and uniqueness of ${\bm{z}}$ are Lipschitz continuity and linear growth conditions for $f$ and $g$ such that if there exist constants $L$ and $C$ such that
$\|f\bigl{(}{\bm{z}}(t_{1});\theta_{f}\bigr{)}-f\bigl{(}{\bm{z}}(t_{2});\theta_{f}\bigr{)}\|$ (11)
$\quad+\|g\bigl{(}{\bm{z}}(t_{1});\theta_{g}\bigr{)}-g\bigl{(}{\bm{z}}(t_{2});\theta_{g}\bigr{)}\|\,\leq\,L\|{\bm{z}}(t_{1})-{\bm{z}}(t_{2})\|$
@@
then a unique strong solution ${\bm{z}}(t)$ exists for all finite $t$. Under these conditions, finite-time blow-up is prevented, ensuring the SDE is well-posed Jia and Benson (2019); Li et al. (2020). Langevin-type SDEs, linear-noise SDEs, and geometric SDEs—variants of Neural SDEs—are also designed to satisfy these conditions, thereby ensuring the existence and uniqueness of their solutions Oh et al. (2024).